OpenStudy (anonymous):

log3(x+1)=2+log3(x-3)

OpenStudy (cwrw238):

first isolate the log terms on one side of the equation then apply laws of logs and convert to exponential form

OpenStudy (cwrw238):

log3(x+1)=2+log3(x-3) log3(x +1) - log3(x-3) = 2 log3[ (x+1)(x-3)] = 2 now convert to exponential form

OpenStudy (anonymous):

log3(x-2)=2?

OpenStudy (cwrw238):

no use the definition of loge i loga x = b then a^b = x

OpenStudy (cwrw238):

in your problem a=3 , b = 2 and x = (x+!)/(x-3)